OpenStudy (anonymous):

Geometry math question. HELP, I need the answer by today? You have a scoop of ice cream 3 inches in diameter sitting on a cone with a 1 inch radius and a height of 12 inches... Assuming all the ice cream melts the cone, determine whether it will fit in the cone or overflow. State the difference of the amount of overflow or the remaining room left in the cone after all the ice cream melts.

OpenStudy (anonymous):

plz help

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

You need the folrmulas for the volume of a sphere and the volume of a cone.

OpenStudy (anonymous):

can u help me solve it plz i dont know what that is

OpenStudy (mathstudent55):

\( \large{V_{sphere} = \dfrac{4}{3} \pi r^3 } \) \( \large {V_{cone} = \dfrac{1}{3} \pi r^2 h } \)

OpenStudy (anonymous):

is this right?

OpenStudy (mathstudent55):

For the sphere (the ice cream scoop), the diameter is 3 in., so the radius is 1.5 in.

OpenStudy (anonymous):

is the picture right ?

OpenStudy (anonymous):

so will that be the correct answer?

OpenStudy (mathstudent55):

That is excellent. A+

OpenStudy (mathstudent55):

Yes

OpenStudy (anonymous):

tank you

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

yw